# “My life is miserable, have to sign 500 autographs everyday”: Exposing Humblebragging, the Brags in Disguise

### v1

@@ 6.2 Qualitative Analysis
T6 is a rare case where our assumption that humblebrags are not present in the datasets used to create our negative samples was violated. After encountering this example, we reviewed our dataset again to ensure no other such cases exist.
@@
T7 and T8 represent classic cases of annotator bias, where annotators attempt to imagine a non-existent context and incorrectly classify the samples as belonging to the positive class. This bias arises because annotators, subconsciously influenced by their task, oversearch for humblebrags in the data they are tasked to annotate for humblebragging detection.
@@ 7 Conclusion and Future Work
We introduced the task of automatic humblebragging detection, formalized through our proposed 4-tuple definition. We benchmarked various machine learning, deep learning, and large language models on this task, providing a comparative analysis against human performance. We also demonstrate that our 4-tuple definition significantly improves the zero-shot capabilities of all decoder models. Additionally, we released a synthetic dataset, HB24, generated using GPT-4o, to facilitate further research. Our experiments and analysis reveal that detecting humblebragging is a challenging task, even for humans. This study lays the groundwork for exploring this intricate linguistic phenomenon and its integration into natural language understanding systems.
Future research could focus on developing improved models for classifying humblebragging, leading to a deeper understanding of this unique communication style. Another promising direction would be the generation of humblebrag captions for images. Lastly, machines could be trained to rewrite direct brags as humblebrags.
@@
The inherent subjectivity of humblebragging complicates the creation of universally agreed-upon labels, as even humans often struggle to classify such statements consistently. Additionally, while machine-generated texts are sophisticated and well-structured, they often lack the spontaneity and imperfections typical of human-authored texts. For instance, the model’s inability to use certain casual or curse words, as well as elongated words like sooooo or goood, which are often present in human-written humblebrags. This creates a mismatch when using synthetic datasets like HB24, which, despite being a valuable resource, may fail to fully capture the linguistic diversity and subtleties of real-world humblebragging, thereby limiting the generalizability of trained models. Moreover, the task itself remains underexplored, with no prior benchmarks or resources, making it difficult to contextualize results within the larger field of natural language understanding.
Appendix A Binary Classification
Text Encoding Generation
$\mathbf{e}=f_{\text{encoder}}(x)$ (2)
where:
  • $x$: Input text.
Binary Classification
The Binary Cross-Entropy (BCE) loss for this task is defined as:
where:
  • $N$: Total number of samples in the dataset.
Objective
Appendix B Data Generation Prompts
@@
Appendix C Pre-Processing
For machine learning classifiers, we incorporated existing pre-processing techniques in two distinct phases. In the first phase, we replaced all emojis with their corresponding verbal explanations, as suggested by Singh et al. (2019), to retain the semantic information. In the second phase, we utilized ekphrasis Baziotis et al. (2017), a specialized text pre-processing tool, to handle hashtags, elongated and repeated words, URLs, and numeric information. The output was further preprocessed by removing stop words and punctuation, constructing unigram and bigram tokens, limiting the vocabulary to the top 10,000 tokens by frequency, and requiring each token to appear in at least two training samples. The resulting numerical representations, created using TF-IDF, were used as inputs for machine learning classifiers such as Support Vector Machine and Logistic Regression.
Appendix D Inference Prompts
@@ 4. Mask Type:
Random Gibberish Prompt
@@ 4. Maple Entanglement:
Appendix E Hyperparameters
The hyperparameters for the experiments are as follows.
For encoder models, BERT-Large-Uncased (340M) was trained with a learning rate of 5e-3, batch size of 16, and 4 epochs, while RoBERTa-Large (355M) used a learning rate of 5e-4, batch size of 32, and 5 epochs. Both models shared a maximum sequence length of 128, a weight decay of 0.01, a warmup ratio of 0.1, and gradient clipping at 1.0.
For decoder models, fine-tuning was conducted using LoRA (Low-Rank Adaptation) with a scaling factor (lora_alpha) of 8, rank (r) of 16, and targeted attention modules (q_proj and k_proj). The dataset was split into 80% training and 20% validation with a random seed of 42. Training employed a cosine learning rate scheduler with a learning rate of 1e-5, weight decay of 0.01, and 10 warmup steps, over 4 epochs with a batch size of 8 and gradient accumulation steps of 4. The maximum sequence length was set to 512 tokens, and the SFTTrainer was utilized for efficient fine-tuning.
Appendix F Confusion Matrices
@@
Appendix G Humblebrag Categories
Category Sezer et al. (2018) Theme Wittels (2012)
Looks and Attractiveness Ugh, Being Hot Sure Can Be Annoying!
Ugh, It’s Tough Being a Model
Ugh, I’m Too Skinny!
Ugh, People Keep Hitting on Me!
Achievements Ugh, Can You Believe They Included Me on This List?
Ugh, I Can’t Believe I Won an Award
Ugh, I’m So Successful
Ugh, I Can’t Believe I Was Mentioned in This Thing!
Performance at Work Ugh, How’d I Get Here??? How Is This My Life???
Money and Wealth Ugh, I Hate Having All This Money!
Intelligence Ugh, I’m a Genius
Personality Ugh, I’m So Humble!
Ugh, It’s Hard Being So Charitable!
Social Life Ugh, It’s So Weird Getting Recognized!
Ugh, I Hate People Wanting My Picture and Autograph All the Time
Ugh, I’m at an Exclusive Event!
Ugh, Being an Author Is Hard!
Miscellaneous Ugh, I Travel Too Much!

### v2

@@ 7.2 Qualitative Analysis
T6 is a rare case where our assumption that humblebrags are not present in the datasets used to create our negative samples was violated. After encountering this example, we reviewed our dataset again to ensure no other such case exists.
@@
T7 and T8 represent classic cases of annotator bias, where annotators attempt to imagine a non-existent context and incorrectly classify the samples as belonging to the positive class. This bias arises because annotators, subconsciously influenced by their task, oversearch for humblebrags in the data as they are tasked to annotate for humblebragging detection.
While our qualitative analysis highlights the nuanced and often subjective nature of humblebrag detection, it also reveals patterns in how models and humans interpret brag-masking cues utilizing our 4-tuple framework. In the next subsection, we compare our framework with another plausible alternative to further reaffirm its suitability for the task.
7.3 4-tuple vs Sentiment Opposition Model
We pit our 4-tuple framework against the Sentiment Opposition Model (SOM) (Appendix N), which is based on detecting incongruity between surface sentiment and intended sentiment. For instance when a statement appears negative or modest but actually conveys an underlying brag.
Table 6 indicates that the 4-tuple Definition (Z+D) outperforms the Sentiment Opposition Model (Z+SOM) in detecting humblebragging, achieving a higher recall and F1-score while maintaining comparable precision. This suggests that the 4-tuple framework more effectively captures both the brag and its masking component, leading to better overall detection. While SOM improves precision by reducing false positives, it sacrifices recall, making it less sensitive to subtle humblebraggings. The higher F1-score of the 4-tuple model confirms its better overall balance between precision and recall, making it a more robust approach for humblebragging detection.
Model A P R F1
Z+D 0.68 0.62 0.88 0.72
Z+SOM 0.66 0.64 0.74 0.68
Table 6: Comparison of Llama-3.1-8B-Instruct using the 4-tuple Definition (Z+D) vs. the Sentiment Opposition Model (SOM). A: Accuracy, P: Precision, R: Recall, F1: F1-score.
7.4 Humblebragging Component Identification
We go a step further in testing the ability of models in identifying the brag and mask components in a given humblebrag by applying our definition. This structured interpretation of humblebragging text demonstrates the practical usefulness of our definition in capturing the nuances of humblebragging enabling models to distinguish the underlying brag from its masked presentation, making detection more accurate and interpretable. Appendix O showcases how different language models process humblebragging using this framework, highlighting variations in their ability to correctly segment and classify such statements. This analysis reinforces the effectiveness of our definition in improving both automated detection and human understanding of humblebragging in natural language.
7.5 Impact of Humblebragging Detection on Downstream Applications
To evaluate the utility of humblebragging detection in a downstream task, we conducted an intended-polarity classification experiment. From the HB-24 test set, we filtered out irony and neutral cases and defined gold labels as follows: humblebragging and bragging were labeled positive, while sarcasm and complaints were labeled negative.
As a baseline, we used RoBERTa-large fine-tuned on the SST-2 dataset Socher et al. (2013), referred to as R-SST2. We then introduced a GPT-based classifier to detect humblebragging and sarcasm, adjusting the sentiment scores accordingly: +1 for humblebrags (to reflect their underlying positivity) and –1 for sarcasm (to correct for overstated positivity). The modified classification module, which integrates these adjustments into the R-SST predictions, is referred to as R-HBSC (HumbleBragging and SarCasm).
Model A P R F1
R-SST2 0.53 0.69 0.53 0.51
R-HBSC 0.82 0.86 0.82 0.83
Table 7: Comparison of vanilla RoBERTa-SST2 (R-SST2) with RoBerta-HBSC (R-HBSC). A: Accuracy, P: Precision, R: Recall, F1: F1-score.
Table 7 shows that adding this pragmatic layer significantly improves sentiment classification across all metrics. This demonstrates that accounting for implicit cues like humblebragging and sarcasm better aligns model predictions with intended sentiment.
@@ 8 Conclusion and Future Work
We introduced the task of automatic humblebragging detection, formalized through our proposed 4-tuple definition. We benchmarked various machine learning, deep learning, and large language models on this task, providing a comparative analysis against human performance. We also demonstrate that our 4-tuple definition significantly improves the zero-shot capabilities of all decoder models. Additionally, we released a synthetic dataset, HB-24, generated using GPT-4o, to facilitate further research. Our experiments and analysis reveal that detecting humblebragging is a challenging task, even for humans. This study lays the groundwork for exploring this intricate linguistic phenomenon and its integration into natural language understanding systems.
Future research could aim to enhance models for identifying humblebragging, fostering a deeper comprehension of this distinct communication style. This may include methods for utilizing contextual cues more effectively. Another valuable direction could involve generating humblebrag captions for images. Additionally, machines could be trained to transform direct brags into humblebrags.
@@
The inherent subjectivity of humblebragging complicates the creation of universally agreed-upon labels, as even humans often struggle to classify such statements consistently. Additionally, while machine-generated texts are sophisticated and well-structured, they often lack the spontaneity and imperfections typical of human-authored texts. For instance, the model’s inability to use certain casual or curse words, as well as elongated words like sooooo or goood, which are often present in human-written humblebrags. This creates a mismatch when using synthetic datasets like HB-24, which, despite being a valuable resource, may fail to fully capture the linguistic diversity and subtleties of real-world humblebragging, thereby limiting the generalizability of trained models. Moreover, the task itself remains underexplored, with no prior benchmarks or resources, making it difficult to contextualize results within the larger field of natural language understanding.
Appendix A On the Prevalence of Humblebragging
Although less common compared to sarcasm or irony, humblebragging is not nonexistent, as evidenced by the Google Trends graph reproduced in Figure 4.
Figure 4: Polpularity of humblebragging.
The graph clearly shows that the popularity of the term has been increasing gradually over time.
Moreover Reddit has a dedicated r/humblebrag subreddit with 188 thousand members. In contrast r/sarcasm has only 39 thousand members while r/irony has only 60 thousand.
In case of Twitter/X, tweetbinder.com returns the following tweet counts over the past week (search performed on 14 Feb 2025):
  • #humblebrag: 72
  • #sarcasm: 200
  • #irony: 200
Note here that 200 is the limit for the free tier searches.
All of the above statistics correspond to the explicit mentions of the phenomena in question. We suspect implicit humblebragging is much more common in social media. But providing statistics about it is all the more difficult.
Additionally, as observed by Wittels in his book Wittels (2012), the phenomena of humblebragging is not recent. In fact only the term for the phenomena is recent. People have been engaging in humblebragging from historical times knowingly or unknowingly.
Appendix B How Humblebragging Differs from Irony, Sarcasm, Bragging and Complaint
While irony, sarcasm, and humblebragging rely on indirect communication, bragging and complaints are more direct. Irony presents a contrast between expectation and reality, often meaning the opposite of what is stated. Sarcasm builds on irony but with a sharper, mocking tone, where the surface meaning appears positive, but the intent is negative. Humblebragging disguises self-promotion within a complaint or self-deprecating remark, appearing negative while aiming to impress.
In contrast, bragging and complaints do not rely on hidden meanings. Bragging openly expresses pride, maintaining a positive tone in both surface meaning and intent. Complaints directly convey dissatisfaction, with both their surface and intended polarity being negative. These distinctions are summarized in Table 8.
Phenomenon Surface Polarity Intended Polarity
Irony Positive or Negative Opposite of surface polarity
Sarcasm Positive Negative
Humblebragging Negative Positive
Bragging Positive Positive
Complaint Negative Negative
Table 8: Surface and Intended Polarity of Different Phenomena
Appendix C Humblebrag Categories
The mapping between Wittel’s humblebrag themes to Sezer’s humblebrag categories is shown in Table 9.
Category Sezer et al. (2018) Theme Wittels (2012)
Looks and Attractiveness Ugh, Being Hot Sure Can Be Annoying!
Ugh, It’s Tough Being a Model
Ugh, I’m Too Skinny!
Ugh, People Keep Hitting on Me!
Achievements Ugh, Can You Believe They Included Me on This List?
Ugh, I Can’t Believe I Won an Award
Performance at Work Ugh, I’m So Successful
Money and Wealth Ugh, I Hate Having All This Money!
Intelligence Ugh, I’m a Genius
Personality Ugh, I’m So Humble!
Ugh, It’s Hard Being So Charitable!
Social Life Ugh, It’s So Weird Getting Recognized!
Ugh, I Hate People Wanting My Picture and Autograph All the Time
Ugh, I’m at an Exclusive Event!
Ugh, Being an Author Is Hard!
Ugh, How’d I Get Here??? How Is This My Life???
Ugh, I Travel Too Much!
Miscellaneous Ugh, I Can’t Believe I Was Mentioned in This Thing!
Appendix D Data Generation Prompts
@@
Appendix E Dataset Quality Assurance
After synthetic data generation, a manual verification was performed by the first two authors of the paper. In this manual verification step, the main aim was to ensure the selection of high-quality samples for the dataset by:
  • Removal of near duplicates.
  • Inclusion of diverse categories of humblebrags that represent real-world scenarios.
Moreover, the 4-tuple definition of humblebragging guided the entire manual filtering stage. For instance, out of the following samples, only one was selected, and the rest were discarded:
  • "Why can’t they just serve normal snacks in first class? Caviar and champagne get so repetitive."
  • "Why do they always offer turn-down service on long-haul flights? Sometimes I just want to make my own bed."
  • "Why does the in-flight chef always insist on preparing gourmet meals? Sometimes I just crave a simple sandwich."
  • "Why do first-class cabins have private suites? I kind of miss the open seating vibe of economy."
This manual filtering stage was followed by a discussion round wherein the two authors discussed both the filtered-in and filtered-out samples. For cases of disagreement, another round of filtering followed by discussion was performed.
Appendix F Binary Classification Framework for Humblebragging Detection
Text Encoding Generation
$\mathbf{e}=f_{\text{encoder}}(x)$ (2)
where:
  • $x$: Input text.
Binary Classification
The Binary Cross-Entropy (BCE) loss for this task is defined as:
where:
  • $N$: Total number of samples in the dataset.
Objective
Appendix G Inference Prompts
@@ 4. Mask Type:
Appendix H Detailed Experimental Setup
Below we outline the hyperparameter settings used for various models along with a brief discussion of the human annotation.
H.1 Machine Learning Classifiers
Thus, for machine learning classifiers, we incorporated existing pre-processing techniques in two distinct phases. In the first phase, we replaced all emojis with their corresponding verbal explanations, as suggested by Singh et al. (2019), to retain the semantic information. In the second phase, we utilized ekphrasis Baziotis et al. (2017), a specialized text pre-processing tool, to handle hashtags, elongated and repeated words, URLs, and numeric information. The output was further pre-processed by removing stop words and punctuation, constructing unigram and bigram tokens, limiting the vocabulary to the top 10,000 tokens by frequency, and requiring each token to appear in at least two training samples. The resulting numerical representations, created using TF-IDF, were used as inputs for machine learning classifiers- Support Vector Machine and Logistic Regression.
H.2 Encoder Models
BERT-Large-Uncased (340M) was trained with a learning rate of 5e-3, batch size of 16, and 4 epochs, while RoBERTa-Large (355M) used a learning rate of 5e-4, batch size of 32, and 5 epochs. Both models shared a maximum sequence length of 128, a weight decay of 0.01, a warmup ratio of 0.1, and gradient clipping at 1.0.
H.3 Decoder Models
We evaluated open-source models from Hugging Face—Qwen2.5-7B-Instruct Yang et al. (2024), Llama 3.1-8B-Instruct Dubey et al. (2024), Gemma 1.1 7B (IT) Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), and Vicuna-7B-v1.5 Zheng et al. (2023)—as well as GPT-3.5 and GPT-4o via OpenAI’s API [^4] https://platform.openai.com/docs/overview. Outputs were limited to two tokens, and each model was run five times to record average metrics. Prompt details are in Appendix G.
Fine-tuning was conducted using LoRA (Low-Rank Adaptation) with a scaling factor (lora_alpha) of 8, rank (r) of 16, and targeted attention modules (q_proj and k_proj). The dataset was split into 80% training and 20% validation with a random seed of 42. Training employed a cosine learning rate scheduler with a learning rate of 1e-5, weight decay of 0.01, and 10 warmup steps, over 4 epochs with a batch size of 8 and gradient accumulation steps of 4. The maximum sequence length was set to 512 tokens, and the SFTTrainer was utilized for efficient fine-tuning.
H.4 Human Performance
Appendix I Inconsistency in Human Annotation
On analyzing the misclassifications made by the lowest-performing human annotator, we observed that the primary factor was not the 4-tuple definition itself. Instead, cultural differences between the source of the tweets and the annotator’s background played a significant role, leading to missed contextual cues.
For instance, the following examples were misclassified by the annotator as they failed to identify celebrity status of people mentiond in the example:
  • Sitting next to Penny Marshall at the Lakers Game. #GEEKINGOUT
  • Tonight: private dinner/event by Miles Davis Estate—his 85th birthday with his family, musicians, media (Beyond humbled/honored to be invited).
  • Watching myself on Larry King. Achievement diminished in ad break by catheter commercial.
To address this issue, future studies should consider recruiting annotators from diverse cultural and geographical backgrounds or implementing mechanisms to provide additional contextual information. However, such interventions are beyond the scope of the present study and are proposed as directions for future research.
Appendix J Random Gibberish Prompt
@@ 4. Maple Entanglement:
Appendix K Text Book Definition
A specific type of brag that masks
the boasting part of a statement
in a faux-humble guise. The false
humility allows the offender to
boast about their achievements
without any sense of shame or guilt.
Humblebrags are usually
self-deprecating in nature.
Appendix L Confusion Matrices
The confusion matrices for Qwen, Gemma, Mistral, Vicuna, BERT and RoBERTa are in Figure 5, Figure 6, Figure 7, Figure 8 and Figure 9 (BERT and RoBERTa) respectively.
@@
Appendix M Performance Degradation Post Fine-Tuning
We did a full fine-tuning on Vicuna but this had no effect on its performance. We suspect that Vicuna, being an older generation model, is not able to cope up with the newer models. This is also evident through Vicuna’s position towards the end of the leaderboards in both Chatbot Arena (https://lmarena.ai/?leaderboard) as well as OpenLLM (https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard#/).
We tried full fine-tuning on Mistral but due to resource constraints we could not explore it much. Our conjecture is that due to the small size of our dataset and the nuanced nature of our task, Mistral could have experienced catastrophic forgetting. This type of behavior is also reported by Li et al. (2024).
For mitigation, future research directions could be increasing the dataset size, augmenting synthetic data with human-written data and exploring hyper-parameter tuning.
Appendix N Sentiment-Opposition Model Prompt
The Sentiment Opposition Model (SOM) for
Humblebragging consists of the following
components:
1. Surface Sentiment (SS):
- The apparent emotional tone of
the statement, usually negative
or neutral (modest).
2. Intended Sentiment (IS):
- The actual meaning the speaker
conveys, which is typically
positive and self-promotional.
3. Sentiment Opposition (SO):
- The contrast between SS and IS.
If SS is negative/neutral but IS
implies positive tone,
opposition exists.
4. Humblebrag Classification:
- If SO exists, classify the
statement as a humblebrag.
- If SO doesnt exist, classify
as a non-humblebrag
Now you are about to classify if a given
sentence is a humblebrag or not using
the above definition.
Appendix O Humblebragging Component Identification Results
See Table 10 for component identification performance of three different models.
Example 1:
Input (Gold): I can’t believe they’d give an idiot like me a PhD lol  (Brag Theme: Achievements; Mask Type: Complaint)
GPT-4o: I can’t believe they’d give an idiot like me a PhD lol  (Brag Theme: Achievement; Mask Type: Self-Deprecating)
Llama-3.1-8B: I can’t believe they’d give an idiot like me a PhD  (Brag Theme: Achievements; Mask Type: Modesty)
Mistral-7B: I can’t believe they’d give an idiot like me a PhD  (Brag Theme: Intelligence; Mask Type: Complaining)
Example 2:
Input (Gold): For the 3rd time in 3 years I’ve been asked to speak at Harvard, but I’ve yet to speak at my alma mater. What’s a girl gotta do @MarquetteU?  (Brag Theme: Achievements; Mask Type: Complaint)
GPT-4o: For the 3rd time in 3 years I’ve been asked to speak at Harvard, but I’ve yet to speak at my alma mater. What’s a girl gotta do @MarquetteU?  (Brag Theme: Achievements / Intelligence; Mask Type: Complaining)
Llama-3.1: For the 3rd time in 3 years I’ve been asked to speak at Harvard, but I’ve yet to speak at my alma mater.What’s a girl gotta do @MarquetteU?  (Brag Theme: Achievements; Mask Type: Complaining)
Mistral-7B: For the 3rd time in 3 years I’ve been asked to speak at Harvard, I’ve yet to speak at my alma mater. What’s a girl gotta do @MarquetteU?  (Brag Theme: Achievement; Mask Type: Complaining)
Example 3:
Input (Gold): Will Twitter be available for me in Paris, Milan, or the Maldives? I hope so bc it won’t in Hong Kong or Singapore.  (Brag Theme: Social Life; Mask Type: Complaint)
GPT-4o: Will Twitter be available for me in Paris, Milan, or the Maldives? I hope so bc it won’t in Hong Kong or Singapore.  (Brag Theme: Social Life / Wealth; Mask Type: Complaining)
Llama-3.1: Will Twitter be available for me in Paris, Milan, or the Maldives? I hope so bc it won’t in Hong Kong or Singapore.  (Brag Theme: Social Life; Mask Type: Modest)
Mistral-7B: Will Twitter be available for me in Paris, Milan, or the Maldives? I hope so bc it won’t in Hong Kong or Singapore.  (Brag Theme: Technology and Connectivity; Mask Type: Complaining)
Table 10: Humblebragging component identification: Model responses with identified brag (red), identified mask (blue), brag theme and mask type.